### v1

@@ 6 Bounds for Reward-Bounded Reachability
Next, we allow $\rho(s)=0$. As the reward of the goal state $g$ does not matter for our purpose, we can w.l.o.g. assume $\rho(g)>0$. The transformation from [12] is not applicable, as it would require division by 00. To circumvent this issue we first construct from $\mathcal{M}$ a CTMC $\mathcal{M}_{>0}$ without states with $\rho(s)=0$. Intuitively, this is done by redirecting the incoming transitions of states with reward 00 to their successors, allowing us to remove any such state from the chain.
@@ 7 Conclusion and Future Work
We introduced $(\varepsilon,\delta)$-bisimulations, a novel type of approximate probabilistic bisimulation for CTMCs. In contrast to notions from the literature such as quasi-lumpability [19, 29, 30], the separate bounds on changes in transition probabilities and changes in exit rates used in $(\varepsilon,\delta)$-bisimulation result in a more flexible, fine-grained notion of behavioral similarity. We established fundamental properties of the notion, and analyzed the difference in timed reachability probabilities of $(\varepsilon,\delta)$-bisimilar states.
We obtained bounds on this difference by uniformizing the chain and applying known results from [18] for the discrete-time setting. Although tight if $\delta=0$, the bounds usually do not perform well if $\delta>0$. Hence, we investigated the special case of $(0,\delta)$-bisimilarity in more detail—also as a stepping stone towards a refined treatment of $(\varepsilon,\delta)$-bisimilarity in future work. For $(0,\delta)$-bisimilar states, we provided bounds based on the error in Erlang CTMCs $\mathcal{E}_{n}$, as well as based on a spectral analysis of the probability matrix $\mathbf{P}$ of the CTMC in question. Furthermore, we extended our results to reward-bounded reachability probabilities, provided that all rewards in the model are nonnegative.
From a theoretical point of view, the behavior of the bounds obtained from the spectral analysis of $\mathbf{P}$ is promising. They are tight and, as can be observed in Figure 7, seem to evolve similar to the actual error. The bounds are, however, stretched by a (large) factor. This is due to the fact that, to obtain the constants $C$, the triangle inequality and maxima are used, causing them to become significantly larger than necessary. This can lead to poor performance of the bounds, in particular if $t$ is small, and necessitates the search for more suitable constants $C^{\prime}\ll C$ in future work. From a practical point of view, applicability of the bounds seems questionable. They tend to over-approximate the error and, depending on the second largest eigenvalue of $\mathbf{P}$, can take a long time (i.e., require large values of $t$) to converge to the actual error, even for small chains. Moreover, computing Jordan canonical forms (and diagonalizations) is computationally expensive and numerically unstable [26, 53, 54]. Consequently, the search for more practically relevant bounds is an important direction for future work.

### v2

@@ 6 Bounds for Reward-Bounded Reachability Probabilities
Next, we allow $\rho(s)=0$. As the reward of the goal state $g$ does not matter for our purpose, we can w.l.o.g. assume $\rho(g)>0$. The transformation from [12] is not applicable, as it could require division by 00. To circumvent this issue we first construct from $\mathcal{M}$ a CTMC $\mathcal{M}_{>0}$ without states with $\rho(s)=0$. Intuitively, this is done by redirecting the incoming transitions of states with reward 00 to their successors, allowing us to remove any such state from the chain.
@@ 7 Conclusion and Future Work
We introduced $(\varepsilon,\delta)$-bisimulations, a novel type of approximate probabilistic bisimulation for CTMCs. In contrast to notions from the literature such as quasi-lumpability [19, 29, 30], the separate bounds on changes in transition probabilities and changes in exit rates used in $(\varepsilon,\delta)$-bisimulation result in a more flexible, fine-grained notion of behavioral similarity. We established fundamental properties of the notion, and analyzed the difference in timed reachability probabilities of $(\varepsilon,\delta)$-bisimilar chains.
We obtained bounds on this difference by uniformizing the chain and applying known results from [18] for the discrete-time setting. Although tight if $\delta=0$, the bounds usually do not perform well if $\delta>0$. Hence, we investigated the special case of $(0,\delta)$-bisimilarity in more detail—also as a stepping stone towards a refined treatment of $(\varepsilon,\delta)$-bisimilarity in future work. For $(0,\delta)$-bisimilar chains, we provided bounds based on the error in Erlang CTMCs $\mathcal{E}_{n}$, as well as based on a spectral analysis of the probability matrix $\mathbf{P}$ of the CTMC in question. Furthermore, we extended our results to reward-bounded reachability probabilities, provided that all rewards in the model are nonnegative.
From a theoretical point of view, the behavior of the bounds obtained from the spectral analysis of $\mathbf{P}$ is promising. They are tight and, as can be observed in Figure 7, seem to evolve similar to the actual error. The bounds are, however, stretched by a (large) factor, which happens since, to obtain the constants $C$, the triangle inequality and maxima are used. This causes the constants to become significantly larger than necessary, leading to poor performance of the bounds in particular if $t$ is small. Therefore, searching for smaller constants $C^{\prime}\ll C$ in future work is in order. From a practical point of view, applicability of the bounds seems questionable. They tend to over-approximate the error and, depending on the second largest eigenvalue of $\mathbf{P}$, can take a long time (i.e., require large values of $t$) to converge to the actual error, even for small chains. Moreover, computing Jordan canonical forms (and diagonalizations) is computationally expensive and numerically unstable [26, 53, 54]. Consequently, the search for more practically relevant bounds is an important direction for future work.